OpenStudy (anonymous):

0.25x^2-0.25x-10.5=0

OpenStudy (anonymous):

\[x _{1, 2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } = \frac{ 0.25 \pm \sqrt{0.0625+ 10.5} }{ 0.5 } = 0.5 \pm \sqrt {10.5625} = 0.5 \pm 3.25\] \[x \in \left\{ 3.75, -2.75 \right\}\]